# Strategic Demand-Planning in Wireless Networks: Can Generative-AI Save Spectrum and Energy?

### v1

@@ Abstract
Wireless communications advance hand-in-hand with artificial intelligence (AI), indicating an interconnected advancement where each facilitates and benefits from the other. This synergy is particularly evident in the development of the sixth-generation technology standard for mobile networks (6G), envisioned to be AI-native. Generative-AI (GenAI), a novel technology capable of producing various types of outputs, including text, images, and videos, offers significant potential for wireless communications, with its distinctive features. Traditionally, conventional AI techniques have been employed for predictions, classifications, and optimization, while GenAI has more to offer. This article introduces the concept of strategic demand-planning through demand-labeling, demand-shaping, and demand-rescheduling. Accordingly, GenAI is proposed as a powerful tool to facilitate demand-shaping in wireless networks. More specifically, GenAI is used to compress and convert the content of various kind (e.g., from a higher bandwidth mode to a lower one, such as from a video to text), which subsequently enhances performance of wireless networks in various usage scenarios such as cell-switching, user association and load balancing, interference management, and disaster scenarios management. Therefore, GenAI can serve a function in saving energy and spectrum in wireless networks. With recent advancements in AI, including sophisticated algorithms like large-language-models and the development of more powerful hardware built exclusively for AI tasks, such as AI accelerators, the concept of demand-planning, particularly demand-shaping through GenAI, becomes increasingly relevant. Furthermore, recent efforts to make GenAI accessible on devices, such as user terminals, make the implementation of this concept even more straightforward and feasible.
@@ I Introduction
The growing demand for coverage and capacity, along with the shift towards the sixth-generation technology standard for mobile networks (6G), is driving the development of technologies that ensure immersive, hyper-reliable, and low-latency communication as well as massive connectivity. 6G introduces innovative applications, such as extended reality (XR), holographic communication, tactile and haptic Internet applications, to name a few, which require much higher levels of key performance indicators (KPIs) compared to applications in legacy networks [1]. Emergence of novel use cases, aligned with the ambitious requirement of KPIs in 6G, necessitates employment of cutting-edge technologies, such as artificial intelligence (AI) and advanced network platforms, including non-terrestrial networks (NTN) [2]. As wireless networks become denser and various network tiers integrate vertically, forming vertical heterogeneous networks (VHetNets), energy consumption and resource utilization increase tremendously, requiring robust and efficient algorithms. Given the scale of future wireless networks and the need to optimize various metrics from both user and network perspectives, wireless networks will inevitably experience a shift towards autonomous and self-optimizing infrastructures [1]. This lies in the maturation of self-organization and the advanced utilization of AI technologies in future wireless networks. The revolutionary enhancements in AI are expected to have a noticeable impact on the telecommunications field to create a sustainable, dynamic, and resilient network that can manage various conditions and scenarios efficiently [3].
@@
In recent years, AI has emerged as a promising tool for designing solutions in wireless networks in the form of machine learning algorithms. Despite efforts to develop efficient algorithms using conventional AI and optimization methods as well as increasing available resources (such as power, time, frequency, and hardware) through advanced technologies, including terahertz (THz) communications, and ultra-massive multiple-input multiple-output (umMIMO) antennas, proposed algorithms still fall short in scenarios where a large number of users demand a huge volume of traffic. Consequently, these algorithms are unable to fulfill users demand and achieve the best possible performance [4]. To tackle this challenge, strategic demand-planning can be considered a potential approach in managing the user traffic, and therefore optimizing the resource utilization and energy consumption of the network.
@@
In wireless networks, as seen in Fig. 1, user demand-planning involves demand-labeling, demand-shaping, and demand-rescheduling. Through demand-labeling, user data can be labeled as either critical or non-critical based on its importance and priority. Accordingly, demand-shaping entails compressing user-demanded data or converting it to another format, depending on the network’s resource utilization status. In scenarios of network congestion, where available resources are insufficient to fulfill user’s requirement, the user data that needs to be sent with high priority and can be shaped without meaning change or information loss, can be shaped, and transmitted using fewer allocated resources. On the other hand, user data with lower priority level or requiring precise transmission of bits can be rescheduled to a time slot where available resources are sufficient for reliable transmission.
@@
Recently, there has been a noticeable interest in employing generative-AI (GenAI) models in the telecommunications field. Several studies have been conducted to investigate the potentials of utilizing GenAI in wireless networks, thereby unveiling new technological possibilities towards 6G [5, 6, 7]. Besides the conventional applications of GenAI in traditional communications that guarantee the accurate transmission and reception of the user’s data, GenAI has the potential to play a crucial role in semantic communications[8]. Additionally, GenAI is a powerful tool to facilitate demand-planning in wireless networks and can be utilized at the user terminal or base stations (BSs) to facilitate demand-planning in scenarios where the available resources are insufficient to accommodate the entire demanded traffic. Through demand-shaping, GenAI improves the performance of wireless networks from resource utilization and energy consumption perspectives, i.e., energy and spectrum are saved. Therefore, demand-planning, and particularly demand-shaping through GenAI, can be a high-potential remedy to enhance future ultra-dense wireless networks, making them more sustainable, dynamic, and resilient.
@@ II Demand-Planning with GenAI: Concept and Models
GenAI technology has undergone rapid enhancements, leading to the development of novel generative models. Each GenAI model comprises various components, and has been designed to achieve a specific objective. In this section, we first discuss the basics of GenAI along with the key models. Next, we introduce the concept of demand-planning, explore how to apply demand-planning in wireless networks, and discuss the associated advantages.
@@ II-A GenAI Basics
Two competing neural network models exist in GANs, engaged in a zero-sum game, i.e., if one model gains, the other loses. The first neural network, referred to as generator, generates fake data samples (i.e., synthetic data) based on the training set, and the second one, named discriminator, determines how realistic the fake data seems. These two parts work together to generate new data with the same statistics as the training set. GAN models have been widely used in the wireless networks literature to analyze transmitted data, and to model traffic patterns and user behaviors. This supports wireless networks in estimating network demand and required resources more efficiently. VAEs are another type of GenAI approaches, that are generative models created by combining deep learning and probabilistic graphic models. VAEs are used to learn complex probabilistic distributions from data sets and generate new data samples. Therefore, they are potential algorithms for anomaly detection and data compression. In particular, VAEs can be instrumental in user’s demand-planning in future wireless networks, enabling the optimization of the resource utilization and energy consumption, where the available resources are not adequate to fulfill system level agreements.
@@
DMs are two-stage deep learning models: in the first stage, noise is incrementally added to real data samples until the data is transformed into complete noise, while in the second stage, the model reduces this noise step-by-step, approaching the original data. In this way, the model learns how to generate realistic data from random data by removing the noise. GenAI DMs are essential in noise reduction and traffic prediction in wireless networks. Transformers, on the other hand, capture long-term dependencies and contextual information within sequences, making them particularly effective for data compression and data optimization tasks. They can efficiently encode sequences into compact representations, reducing the amount of data needed to capture important information. Furthermore, compared to conventional sequential models, transformers can process entire sequences in parallel, leading to faster and more efficient optimization processes. The comparisons of GenAI models are summarized in Table I.
In order to justify that the use and integration of GenAI into devices can be regarded as a near-future concept, it is worth noting that major manufacturers, such as Apple, have started forming partnerships with AI companies for on-device GenAI usage [10], thereby proving the applicability of the concept introduced in this work.
@@ II-B1 Demand-labeling
Figure 3: GenAI demand-shaping usage scenarios in wireless networks. The GenAI logo in the middle of the figure has been created by the DALL·E 2 which is a text-to-image model developed by OpenAI.
@@ II-B2 Demand-shaping
With the help of the aforementioned models, GenAI can compress or convert the user data to optimize the network’s traffic through demand-shaping. Data compression can support shaping user data into a lower volume, which requires lower data rates and fewer allocated resources to be transmitted. For instance, the size of high-resolution videos can be reduced through video compression algorithms provided by GenAI at the transmitter, and the shaped video can be converted back to its original format (i.e., decompression) using DMs at the receiver side. Data conversion is another approach to perform demand-shaping in wireless networks. In particular, contents such as multimedia and video, which require relatively more resources, can be converted to a lower-data-rate format before being transmitted through the network. Another way to shape the demand can be by analyzing and optimizing the content of the data. In other words, GenAI can make comprehensions from the meaning of the data, and optimize the content. To this end, voice and text messages can be summarized using GenAI, without loss of inference from its original type, in order to compress the user data to further relax the network traffic.
@@ II-B3 Demand-rescheduling
Fig. 2 illustrates the concept and methodology of demand-planning through GenAI in wireless networks. The process begins with user data labeling and proceeds to demand-shaping. Pre-trained GAN models generate synthetic data to complement the collected data in demand-shaping; therefore, a comprehensive dataset is ensured. The user demand can then be compressed and optimized using VAEs, DMs, or transformer algorithms. Hence, the resource allocation and network performance become more efficient thanks to the optimized dataset as well as the compression process. Additionally, the transmission of non-critical data can be rescheduled when necessary.
@@ III Usage Scenarios in Wireless Networks
Employing GenAI models for demand-planning in wireless networks provides advantages in various usage scenarios. Particularly, demand-planning can be implemented in uplink or downlink communications to improve the resource utilization of the wireless networks, in scenarios where the resources are insufficient to fulfill the users’ demand. According to the network’s available resources and specific requirements of the users, a GenAI model can be employed, at the user terminal and/or the network side, i.e., at the BS, to facilitate demand-planning.
@@ III-A Sustainability through Cell-Switching
Cell-switching is a viable solution that seeks idle or lightly loaded SBSs, offloads their traffic to MBSs, and puts them into sleep mode as they are no longer needed to operate [11]. The operating principle of cell-switching is based on the density of the resource blocks (RBs) of BSs (i.e., data traffic loads). GenAI-based demand-planning allows for adjusting the RBs demanded by each user, thereby optimizing the resource utilization of a BS. Therefore, the available capacity of the MBS increases due to the lower amount of data offloaded from switched-off BSs, as a result of demand-planning process, providing more switching off opportunities. Consequently, the energy consumption of the overall wireless network will be decreased, leading to a more sustainable networking. The impact of demand-shaping through GenAI on the energy consumption of wireless networks is depicted in Fig. 4. For this purpose, the daily data demands of the users from the city of Milan, collected by Telecom Italia [12], are considered and cell-switching is performed with an exhaustive search algorithm in networks with diverse set of BSs (1 MBS, and 2, 4, 6 SBSs). We consider that the actual user demand is compressed by GenAI using various compression ratios: 0%, 20%, 40%, 60%, and 80%. It is observed from the findings in Fig. 4 that by increasing the compression ratio of demand-shaping algorithm, the network’s total energy consumption decreases. In more complex and dense networks, consisting of NTN platforms (e.g., HIBS [11]), demand-shaping will play a substantial role in optimizing energy consumption. For instance, after shaping the demands through GenAI, NTN elements like HIBS can host the offloaded users, increasing more switching-off possibilities that is translated into even more energy saving as HIBS is envisioned to be empowered with green-friendly energy sources.
@@ III-B User Association and Load Balancing
Among these limitations is the restricted wireless backhaul link capacity of aerial platforms, such as HIBS, or the terrestrial BSs lacking fiber facilities. Due to this backhaul link capacity constraint, BSs are limited in the number of users they can support. By implementing GenAI at BSs with wireless backhaul capacity, user data can be compressed to a format that requires a lower data rate, preventing congestion on the wireless backhaul link. To this end, the BS analyzes the users’ demanded traffic, followed by deciding to implement the GenAI algorithm to compress the users’ data based on available capacity in the wireless backhaul link. Considering the type of users’ data, the BS may also decide to select a specific group of users for demand-shaping.
@@ III-C Interference Management
VHetNets enhance the coverage and capacity of terrestrial networks and contribute to making wireless networks more resilient. International Telecommunications Union (ITU) at the World Radio Conference 2023 (WRC-23) took a significant step by allocating additional spectrum, which is already used by IMT, to HIBS [13]. This advancement is crucial as it facilitates the integration of HIBS with the existing terrestrial wireless networks in a harmonized spectrum manner, where various network tiers share the same frequency band. Nevertheless, ultra-dense VHetNets, particularly those operating in a harmonized spectrum, experience performance issues due to propagated inter/intra-tier interference [14]. Consequently, designing effective interference management mechanisms becomes crucial.
One of the most common approaches to optimize interference in wireless networks is through designing robust and efficient resource management schemes[14]. However, despite the efforts towards developing efficient resource allocation algorithms, the continuous rise in traffic demands from mobile users will be an obstacle. This means that in certain scenarios, the interference management algorithms, developed through resource allocation schemes, fail to deliver the expected improvements, specifically due to the limited amount of available resources. Demand-planning through GenAI increases the degree of freedom in allocating resources to users. Employing GenAI demand-planning, BSs can shape the user data and manage traffic more effectively, leading to enhanced interference management. Fig. 5 represents the sum spectral efficiency in a VHetNet, consisting of 1 HIBS and 4 MBSs, where the total available RBs are shared by 50 users. As depicted in Fig. 5, by compressing the demanded data by the users, the RBs can be allocated to the users more efficiently, leading to decreased interference and increased sum spectral efficiency.

### v2

@@ Abstract
Generative-AI (GenAI), a novel technology capable of producing various types of outputs, including text, images, and videos, offers significant potential for wireless communications. This article introduces the concept of strategic demand-planning through demand-labeling, demand-shaping, and demand-rescheduling. Accordingly, GenAI is proposed as a powerful tool to facilitate demand-shaping in wireless networks. More specifically, GenAI is used to compress and convert the content of various types (e.g., from a higher bandwidth mode to a lower one, such as from a video to text), which subsequently enhances performance of wireless networks in various usage scenarios, such as cell-switching, user association and load balancing, interference management, as well as disasters and unusual gatherings. Therefore, GenAI can serve a function in saving energy and spectrum in wireless networks. With recent advancements in AI, including sophisticated algorithms like large language models and the development of more powerful hardware built exclusively for AI tasks, such as AI accelerators, the concept of demand-planning, particularly demand-shaping through GenAI, becomes increasingly relevant. Furthermore, recent efforts to make GenAI accessible on devices, such as user terminals, make the implementation of this concept even more straightforward and feasible.
@@ I Introduction
The growing demand for coverage and capacity, along with the shift towards the sixth-generation technology standard for mobile networks (6G), is driving the development of technologies that ensure immersive, hyper-reliable, and low-latency communication, as well as massive connectivity. Emergence of novel use cases, aligned with the ambitious requirement of key performance indicators (KPIs) in 6G, necessitates employment of cutting-edge technologies, such as artificial intelligence (AI) and advanced network platforms, including non-terrestrial networks (NTN) [1],[2]. As wireless networks become denser and various network tiers integrate vertically, forming vertical heterogeneous networks (VHetNets), energy consumption and resource utilization increase tremendously, requiring robust and efficient algorithms. Given the scale of future wireless networks and the need to optimize various metrics from both user and network perspectives, wireless networks will inevitably experience a shift towards autonomous and self-optimizing infrastructures [1]. This lies in the maturation of self-organization and the advanced utilization of AI technologies in future wireless networks. The revolutionary enhancements in AI are expected to have a noticeable impact on the telecommunications field to create a sustainable, dynamic, and resilient network that can manage various conditions and scenarios efficiently [3].
@@
In recent years, AI has emerged as a promising tool for designing solutions in wireless networks. Despite efforts to develop efficient algorithms using conventional AI and optimization methods, as well as increasing available resources (such as power, time, frequency, and hardware) through advanced technologies, including terahertz (THz) communications, and ultra-massive multiple-input multiple-output (umMIMO) antennas, proposed algorithms still fall short in scenarios where a large number of users demand a huge volume of traffic. Consequently, these algorithms are unable to fulfill users demand and achieve the best possible performance [4]. To tackle this challenge, strategic demand-planning can be considered a potential approach in managing the user traffic, and therefore optimizing the resource utilization and energy consumption.
@@
In wireless networks, as seen in Fig. 1, user demand-planning involves demand-labeling, demand-shaping, and demand-rescheduling. Through demand-labeling, user data can be labeled as either critical or non-critical based on its importance and priority. Accordingly, demand-shaping entails compressing user-demanded data or converting it to another format, depending on the network’s resource utilization status. In scenarios of network congestion, where available resources are insufficient to fulfill user’s requirement, the user data that needs to be sent with high priority and can be shaped without meaning change or information loss, can be shaped, and transmitted using fewer allocated resources. On the other hand, user data with a lower priority level or requiring precise transmission of bits can be rescheduled to a time slot where available resources are sufficient for reliable transmission.
@@
Recently, there has been a noticeable interest in employing generative-AI (GenAI) models in the telecommunications field. To this end, several studies have explored the potential of utilizing GenAI for network planning and optimization [5],[6]. Furthermore, GenAI has been considered a promising tool for unveiling new technological possibilities for 6G [7],[8]. Besides the conventional applications of GenAI in traditional communications that guarantee the accurate transmission and reception of the user’s data, GenAI has the potential to play a crucial role in semantic communications[9]. Additionally, GenAI can be utilized at the user terminal or base stations (BSs) to facilitate demand-planning in wireless networks in scenarios where the available resources are insufficient to accommodate the entire demanded traffic. Through demand-shaping, GenAI improves the performance of wireless networks from resource utilization and energy consumption perspectives, i.e., energy and spectrum are saved. Therefore, demand-planning, and particularly demand-shaping through GenAI, can be a high-potential remedy to enhance future ultra-dense wireless networks, making them more sustainable, dynamic, and resilient.
@@ II Demand-Planning with GenAI: Concept and Models
GenAI technology has undergone rapid enhancements, leading to the development of novel generative models. Each GenAI model comprises various components, and has been designed to achieve a specific objective. In this section, we first discuss the basics of GenAI along with the key models. Next, we introduce the concept of demand-planning, explore its application in wireless networks, and discuss the associated advantages.
@@ II-A GenAI Basics
Two competing neural network models exist in GANs, engaged in a zero-sum game, i.e., if one model gains, the other loses. The first neural network, referred to as generator, generates fake data samples (i.e., synthetic data) based on the training set, and the second one, named discriminator, determines how realistic the fake data seems. These two parts work together to generate new data with the same statistics as the training set. GAN models have been widely used in the wireless networks literature to analyze transmitted data, and to model traffic patterns and user behaviors. This supports wireless networks in estimating network demand and required resources more efficiently. VAEs are another type of GenAI approaches, that are generative models created by combining deep learning and probabilistic graphic models. VAEs are used to learn complex probabilistic distributions from data sets and generate new data samples. Therefore, they are potential algorithms for anomaly detection and data compression. In particular, VAEs can be instrumental in user’s demand-planning in future wireless networks, enabling the optimization of the resource utilization and energy consumption, where the available resources are not sufficient.
@@
DMs are two-stage deep learning models: in the first stage, noise is incrementally added to real data samples until the data is transformed into complete noise, while in the second stage, the model reduces this noise step-by-step, approaching the original data. In this way, the model learns how to generate realistic data from random data by removing the noise. GenAI DMs are essential in noise reduction and traffic prediction in wireless networks. Transformers, on the other hand, capture long-term dependencies and contextual information within sequences, making them particularly effective for data compression and data optimization tasks. Furthermore, compared to conventional sequential models, transformers can process entire sequences in parallel, leading to faster and more efficient optimization processes. It is worth mentioning that large language models (LLMs) are trained and developed using transformer architectures and have potential applications in telecommunications networks [7, 9]. LLMs can optimize data and network performance by prioritizing critical information and adapting to contextual changes. Additionally, they can predict data sequences, enabling the reconstruction of lost data, which ultimately enhances the reliability of communications. The comparisons of GenAI models are summarized in Table I.
It is worth noting that major manufacturers, such as Apple, have started forming partnerships with AI companies for on-device GenAI usage [11], thereby proving the applicability of the concept introduced in this work.
@@ II-B1 Demand-labeling
Figure 3: GenAI demand-shaping usage scenarios in wireless networks. The GenAI logo in the middle of the figure has been created by the DALL·E 2.
@@ II-B2 Demand-shaping
With the help of the aforementioned models, GenAI can compress or convert the user data to optimize the network’s traffic through demand-shaping. Data compression can support shaping user data into a lower volume, which requires lower data rates and fewer allocated resources to be transmitted. For instance, the size of high-resolution videos can be reduced through video compression algorithms provided by GenAI at the transmitter, and the shaped video can be converted back to its original format (i.e., decompression) using DMs at the receiver side. To achieve the optimal performance, the GenAI models at the transmitter and receiver should be consistent. While any compression algorithm can reduce data size, GenAI excels with its content-aware and regenerative capabilities. It analyzes data to identify key sections and applies variable compression based on their importance, compressing less critical parts more heavily. Additionally, it can regenerate content to correct any compression-related errors.
Data conversion is another approach to perform demand-shaping in wireless networks. In particular, contents such as multimedia and video, which require relatively more resources, can be converted to a lower-data-rate format before being transmitted through the network. Another way to shape the demand can be by analyzing and optimizing the content of the data. In other words, GenAI can make comprehensions from the meaning of the data, and optimize the content. To this end, voice and text messages can be summarized using GenAI, without loss of inference from its original type, to compress the user data to further relax the network traffic.
@@ II-B3 Demand-rescheduling
Fig. 2 illustrates the concept and methodology of demand-planning through GenAI in wireless networks. The process begins with user data labeling and proceeds to demand-shaping. Pre-trained GAN models generate synthetic data to complement the collected data in demand-shaping; therefore, a comprehensive dataset is ensured. The user demand can then be compressed and optimized using VAEs, DMs, or transformer algorithms. Hence, the resource allocation and network performance become more efficient thanks to the optimized dataset as well as the compression process. Additionally, the transmission of non-critical data can be rescheduled when necessary. In conclusion, due to the gap in a single GenAI model being able to perform these tasks in the literature, we envision using different GenAI models to perform various tasks. When these models are used together, they can execute the desired demand planning operations.
@@ III Usage Scenarios in Wireless Networks
Employing GenAI models for demand-planning in wireless networks provides advantages in various usage scenarios. Particularly, demand-planning can be implemented in uplink or downlink communications to improve the resource utilization of the wireless networks, in scenarios where the resources are insufficient to fulfill the users’ demand. According to the network’s available resources and specific requirements of the users, a GenAI model can be employed, at the user terminal and/or the network side to facilitate demand-planning.
@@ III-A Sustainability through Cell-Switching
As wireless networks’ energy consumption increases, so does the amount of carbon released into the atmosphere, which is causing a failure to meet climate change commitments, one of the most significant challenges of our time. One important technology for reducing energy consumption in wireless networks is cell switching, a viable solution that seeks idle or lightly loaded SBSs, offloads their traffic to MBSs, and puts them into sleep mode as they are no longer needed to operate [12]. The operating principle of cell-switching is based on the density of the resource blocks (RBs) of BSs (i.e., data traffic loads). GenAI-based demand-planning allows for adjusting the RBs demanded by each user, thereby optimizing the resource utilization of a BS. Hence, the available capacity of the MBS increases due to the lower amount of data offloaded from switched-off BSs, as a result of demand-planning process, providing more switching off opportunities. Consequently, the energy consumption of the overall wireless network will be decreased, leading to a more sustainable networking.
The impact of demand-shaping through GenAI on the energy consumption of wireless networks is depicted in Fig. 4. For this purpose, the daily data demands of the users from the city of Milan, collected by Telecom Italia [13], are considered and cell-switching is performed with an exhaustive search algorithm in networks with diverse set of BSs (1 MBS, and 2, 4, 6 SBSs). We consider that the actual user demand is compressed by GenAI using various compression ratios: 0%, 20%, 40%, 60%, and 80%. It is observed from the findings in Fig. 4 that by increasing the compression ratio of demand-shaping algorithm, the network’s total energy consumption decreases. In more complex and dense networks, consisting of NTN platforms (e.g., HIBS [12]), demand-shaping will play a substantial role in optimizing energy consumption. For instance, after shaping the demands through GenAI, NTN elements like HIBS can host the offloaded users, increasing more switching-off possibilities that is translated into even more energy saving as HIBS is envisioned to be empowered with green-friendly energy sources. An important point to consider is the trade-off between the energy savings achieved through GenAI-assisted cell switching and the increased energy consumption by GenAI algorithms. The increased energy consumption of AI/GenAI implementations presents a challenge in terms of carbon emissions that must be addressed during the design process to ensure that the energy savings significantly outweigh the energy consumed by the AI.
@@ III-B User Association and Load Balancing
Among these limitations is the restricted wireless backhaul link capacity of aerial platforms, such as HIBS, or the terrestrial BSs lacking fiber facilities. Due to the limited available spectrum resources and the backhaul link capacity constraint, BSs are limited in the number of users they can support. By implementing GenAI at BSs with wireless backhaul capacity, user data can be compressed to a format that requires less amount of allocated spectrum, and a lower data rate, preventing congestion on the wireless backhaul link. To this end, the BS analyzes the users’ demanded traffic, followed by deciding to implement the GenAI algorithm to compress the users’ data based on available capacity and spectrum in the wireless backhaul link. Considering the type of users’ data, the BS may also decide to select a specific group of users for demand-shaping.
@@ III-C Interference Management
VHetNets enhance the coverage and capacity of terrestrial networks and contribute to making wireless networks more resilient. Nevertheless, ultra-dense VHetNets, particularly those operating in a harmonized spectrum, where various network tiers share the same frequency band, experience performance issues due to propagated inter/intra-tier interference [14]. Consequently, designing effective interference management mechanisms becomes crucial.
One of the most common approaches to optimize interference in wireless networks is through designing robust and efficient spectrum allocation schemes[14]. However, despite the efforts towards developing efficient spectrum allocation algorithms, the continuous rise in traffic demands from mobile users will be an obstacle. This means that in certain scenarios, the interference management algorithms, developed through spectrum allocation schemes, fail to deliver the expected improvements, specifically due to the limited amount of available spectrum resources. Demand-planning through GenAI increases the degree of freedom in allocating spectrum to users. Employing GenAI demand-planning, BSs can shape the user data and manage traffic more effectively, leading to enhanced interference management. Fig. 5 represents the sum spectral efficiency in a VHetNet, consisting of 1 HIBS and 4 MBSs, where the total available RBs are shared by 50 users. As depicted in Fig. 5, by compressing the demanded data by the users, the RBs can be allocated to the users more efficiently, leading to decreased interference and increased sum spectral efficiency.